Find the x-intercept of a line that passes through (-3,4) and is perpendicular to the graph of x-2y=7.



* All i know is when the slope is perpendicular to another it is the oppopsite. For example,
3/4
the perpendicular slope of this is
-4/3

Answer by b_gulsum(6)

the slope of 1st line is 1/2, since
2y=x-7
y=x/2-7/2.
Therefore, the slope of 2nd line is -2.
m:slope
(x1,y1): (-3,4)
the line we're looking for is:
m=(y-y1)/(x-x1)
-2=(y-4)/(x+3)
y=-2x-2
Then, x-intercept of this line: where y=0, x=-1